# Dynamics of Dehn Twists in the Outer Automorphism Group of a Free Group

### v1

@@ 1. Introduction
Let $\mathbb{F}=\mathbb{F}_{g}$ be a free group of rank $g$. The automorphism group $\operatorname{Aut}(\mathbb{F})$ contains the inner automorphism group $\operatorname{Inn}(\mathbb{F})$ as a normal subgroup, and the quotient
@@
is called the outer automorphism group of $\mathbb{F}$. Given a one-edge $\mathbb{Z}$-splitting $T$ of $\mathbb{F}$, one can associate a Dehn twist (or generalized transvection) $\delta_{\vec{T}}\in\operatorname{Out}(\mathbb{F})$, defined as the outer automorphism induced by twisting along $T$.
@@
The problem of constructing free subgroups of $\operatorname{Out}(\mathbb{F})$ has been extensively studied. Bestvina, Feighn, and Handel [MR1765705] proved that $\operatorname{Out}(\mathbb{F})$ satisfies the Tits alternative: every subgroup is either virtually solvable or contains a free subgroup of rank two. Kapovich and Lustig [MR2652906] showed that sufficiently large powers of two hyperbolic iwip elements generate either a virtually cyclic subgroup or a nonabelian free group. Uyanik [MR3954286] obtained a related result in the setting of hyperbolic extensions. Clay and Pettet [MR2720227] proved that if two one-edge $\mathbb{Z}$-splittings are filling, then sufficiently large powers of the associated Dehn twists generate a free group of rank two consisting of atoroidal fully irreducible elements except for the conjugacy classes of the twists themselves. Gültepe [MR3677931] provided another criterion under which powers of two Dehn twists generate a free subgroup of rank two.
A complementary development is due to Taylor [MR3343354], who proved that every right-angled Artin group embeds quasi-isometrically into some $\operatorname{Out}(\mathbb{F})$. He constructed a suitable collection of fully irreducible outer automorphisms supported on free factors, whose sufficiently large powers generate the desired RAAG. Since these automorphisms are fully irreducible, Dehn twists do not appear in his construction. Our work may therefore be viewed as a counterpart to Taylor’s theorem: we construct right-angled Artin subgroups of $\operatorname{Out}(\mathbb{F})$ generated instead by powers of Dehn twists.
Theorem 1.1 gives a broad and explicit method for constructing RAAG subgroups inside $\operatorname{Out}(\mathbb{F})$. However, the compatibility condition does not encompass all right-angled Artin phenomena arising from Dehn twists. For instance, $\operatorname{Out}(\mathbb{F}_{g})$ contains a direct product of 2g42g-4 free groups generated by Dehn twists (Horbez and Wade [MR4069231], and Bridson and Wade [MR4792836]) and a free-abelian subgroup of rank 2g32g-3 generated by Dehn twists (Feighn and Handel [MR2496054]), yet the associated one-edge $\mathbb{Z}$–splittings in these examples are not compatible. Thus, Theorem 1.1 captures a large and well-structured class of RAAGs generated by powers of Dehn twists, but it does not apply to every such instance found in $\operatorname{Out}(\mathbb{F}_{g})$.
We approach the proof of Theorem 1.1 by working in the mapping class group of the doubled handlebody $\mathbb{M}=\mathbb{M}_{g}=\#_{g}(\mathbb{S}^{1}\times\mathbb{S}^{2})$. Laudenbach [MR356056] established the following short exact sequence:
@@
Brendle, Broaddus, and Putman [MR4557874] later proved that this sequence splits; in particular, there exists a section $\operatorname{Out}(\mathbb{F})\to\operatorname{MCG}(\mathbb{M})$. In this setting, a Dehn twist may be interpreted as an isotopy class of a homeomorphism twisting $\mathbb{M}$ along an essential torus. Gültepe [MR3677931] likewise studied Dehn twists realized as twists along essential tori.
@@
One can formulate analogous criteria in the cases of two essential spheres or of an essential sphere together with an essential torus. In addition, one obtains a bigon criterion in this setting: an essential simple closed curve and an essential sphere (or an essential torus) intersect minimally if and only if they do not form a bigon. We note that Hatcher [MR1314940] introduced a normal form for sphere systems, and this normal form guarantees that any two sphere systems placed in it are in minimal position.
Meanwhile, Theorem 1.1 is proved by a ping–pong argument on the set of homotopy classes of essential simple closed curves. We employ a variation of the ping–pong method due to Koberda [MR3000498]. Careful control of intersection numbers within this framework yields the proof of Theorem 1.1.
In addition, we construct a compact topological space $\overline{(p\circ i_{*})(\mathcal{C})}$ on which $\operatorname{Out}(\mathbb{F})$ acts continuously. This space contains a dense subset consisting of points represented by homotopy classes of essential simple closed curves. The definition of $\overline{(p\circ i_{*})(\mathcal{C})}$ is closely analogous to that of the projective space $\mathcal{PML}(S)$ of measured laminations on a surface $S$. In $\mathcal{PML}(S)$, the action of a Dehn twist is parabolic, and a Dehn twist in $\operatorname{Out}(\mathbb{F})$ exhibits the same dynamical behavior on $\overline{(p\circ i_{*})(\mathcal{C})}$, as described below.
@@
Let $T$ be a homotopy class of an essential torus with the core $\gamma$. Let $\alpha$ be a homotopy class of an essential simple closed curve. Then, we have $i(\alpha,T)>0$ if and only if $\delta_{\vec{T}}^{n}(\alpha)$ converges to $\gamma$ as $n\to\infty$.
1.1. Acknowledgment
The author thanks Sang-hyun for many helpful and inspiring conversations during the preparation of this work.
@@ 2.1. Dehn twist automorphisms of free groups
An $\mathbb{F}$-tree $T$ is called a one-edge $\mathbb{Z}$-splitting if the action of $\mathbb{F}$ on the set of edges is transitive and a stabilizer of an edge is an infinite cyclic subgroup of $\mathbb{F}$. Let $\vec{E}T$ be the set of orientations of edges of $T$. Then there exists a orientation reversing $\vec{E}T\to\vec{E}T$, that is, $(\vec{e})$ is the reverse orientation of $\vec{e}$ for each $\vec{e}\in\vec{E}T$. Let $o:\vec{E}T\to\mathbb{F}$ be a map defined by
@@ 2.2. Dehn twists of doubled handlebodies
Let $\operatorname{MCG}(\mathbb{M})$ denote the mapping class group of $\mathbb{M}$, that is, the group of isotopy classes of orientation-preserving homeomorphisms. Laudenbach [MR356056] shows the following short exact sequence holds:
@@
Brendle, Broaddus and Putman [MR4557874] discover that this sequence is splitting, that is, there exists a section $\operatorname{Out}(\mathbb{F})\hookrightarrow\operatorname{MCG}(\mathbb{M})$. It is therefore natural to refer to the image of a Dehn twist in $\operatorname{MCG}(\mathbb{M})$ simply as a Dehn twist. What kind of a homeomorphism is associated with a Dehn twist? To answer to this question, we need to dive into the geometry of $\mathbb{M}$.
@@
Note if two essential tori are homotopic, then they give the same Dehn twist as outer automorphisms. By Brendle, Broaddus and Putman [MR4557874], this outer automorphism can be considered as a mapping class. Note the mapping class group consists of isotopy classes of orientation-preserving homeomorphisms. So the following holds.
Lemma 2.7.
@@
We will abuse notation by writing $T$ for both a homotopy class of an essential torus $t$ and the associated dual tree of $t$. Similarly, we use $\vec{T}$ to denote either of their oriented versions. With this convention, the Dehn twist along $\vec{T}$ will be denoted by $\delta_{\vec{T}}$.
@@ 3.1. Bigon, bihedron and minimal position
The following proof is adapted from Farb and Margalit [MR2850125].
@@
    Suppose there are disks $D\subset t$ and $D^{\prime}\subset t^{\prime}$ such that $D\cup D^{\prime}$ is a sphere that bounds a $3$-ball. (We say $t$ and $t^{\prime}$ form a bihedron.) Then one can homotope $t$ by locating $D$ across $D^{\prime}$, which removes a circle $D\cap D^{\prime}$. Then the number of circles of $t\cap t^{\prime}$ decreases.
@@
    Suppose there are pairs of pants $P\subset t$ and $P^{\prime}\subset t^{\prime}$ such that $P\cup P^{\prime}$ is a twice-punctured torus whose hole bounds a disk. (We say $t$ and $t^{\prime}$ form a bigon.) Then one can homotope $t$ by enlarging $P$ so that $P\cap P^{\prime}$ becomes a circle. Then the number of circles of $t\cap t^{\prime}$ decreases.
@@
Let $s$ and $t$ be essential spheres, essential tori, or one of each. Then, the number of circles of $s\cap t$ is exactly $i([s],[t])$ if and only if $s$ and $t$ form neither a bigon nor a bihedron.
@@
Choose two such circles, say $C_{1}$ and $C_{2}$. Then there exist arcs $a\subset\hat{s}\times\{0\}$ and $b\subset S$, each joining $C_{1}$ to $C_{2}$. Let $P$ be a regular neighborhood of $H(C_{1}\cup a\cup C_{2})$ in $s$, and let $Q$ be a regular neighborhood of $H(C_{1}\cup b\cup C_{2})$ in $t$. Then $P$ and $Q$ are pairs of pants that together form a bigon.
@@
By Laudenbach [MR356056], two homotopic sphere systems are isotopic. Hatcher [MR1314940] show that two sphere systems can be in “normal form.” This is equivalent to minimal position. That is, two sphere systems are in normal form if and only if they are in minimal position. Refer to Horbez’s article [MR3020214].
The situation is not different if we consider an essential torus and a sphere system. Gültepe [MR3043125] defines the normal form of a torus with respect to a sphere system. This is equivalent to minimal position.
@@ 3.2. Intersection number and Dehn twist
Let $T$ and $T^{\prime}$ be homotopy classes of essential tori. If $i(\delta_{\vec{T}}(T^{\prime}),T^{\prime})=0$ for some orientation of $T$, then we have $i(T,T^{\prime})=0$.
@@
Suppose $i(T,T^{\prime})>0$, and choose representatives $t\in T$ and $t^{\prime}\in T^{\prime}$ in minimal position. Let $N(t)$ be a regular neighborhood of $t$, together with a homeomorphism
$\phi:N(t)\to t\times[0,1].$
Let us assume that $N(t)$ is small enough that $\phi\big(N(t)\cap t^{\prime}\big)$ is a disjoint union of annuli. Let $\delta_{\vec{t}}$ be a representative of $\delta_{\vec{T}}$ whose support is contained in $N(t)$. Since $t$ and $t^{\prime}$ intersect, it follows that $\delta_{\vec{t}}(t^{\prime})$ and $t^{\prime}$ also intersect. Thus it remains to show that $\delta_{\vec{t}}(t^{\prime})$ and $t^{\prime}$ are in minimal position.
Suppose $\delta_{\vec{t}}(t^{\prime})$ and $t^{\prime}$ form a bihedron. Then there exist disks $D_{1}\subset\delta_{\vec{t}}(t^{\prime})$ and $D_{2}\subset t^{\prime}$ such that $D_{1}\cup D_{2}$ is a sphere bounding a 3-ball. Note that $D_{1}\cap D_{2}\subset N(t)\cap t^{\prime}$. By enlarging $D_{2}$, we may assume its boundary lies in $t\cap t^{\prime}$. And homotope $D_{1}$ such that the interior of $D_{1}$ is disjoint from $t^{\prime}$. Let $\psi:t\times[0,1]\to t$ denote the projection. Then $\psi\phi(D_{1})$ is a disk whose boundary is contained in $t\cap t^{\prime}$. Thus $\psi\phi(D_{1})$ together with $D_{2}$ forms a bihedron in $t\cup t^{\prime}$, contradicting the assumption that $t$ and $t^{\prime}$ are in minimal position.
Suppose $\delta_{\vec{t}}(t^{\prime})$ and $t^{\prime}$ form a bigon. Then there exist pairs of pants $P_{1}\subset\delta_{\vec{t}}(t^{\prime})$ and $P_{2}\subset t^{\prime}$ such that $P_{1}\cup P_{2}$ is a twice-punctured torus bounding a disk. By enlarging $P_{2}$, we may assume its two boundary components lie in $t\cap t^{\prime}$. On the other hand, $\psi\phi(P_{1})$ is a pair of pants contained in $t$. So $\psi\phi(P_{1})$ and $P_{2}$ form a bigon in $t\cup t^{\prime}$, again contradicting the minimal position of $t$ and $t^{\prime}$.
Since both cases lead to contradictions, $\delta_{\vec{t}}(t^{\prime})$ and $t^{\prime}$ must be in minimal position. As they intersect, we conclude that
$i(\delta_{\vec{T}}(T^{\prime}),T^{\prime})>0,$
as desired. ∎

### v2

@@ 1. Introduction
Let $\mathbb{F}=\mathbb{F}_{g}$ be a free group of rank $g$. An $\mathbb{F}$-tree is a simplicial tree on which $\mathbb{F}$ acts simplicially without edge inversions. An $\mathbb{F}$-tree is called a $\mathbb{Z}$-splitting if every edge stabilizer is infinite cyclic. It is said to be one-edge if the induced action of $\mathbb{F}$ on the set of edges is transitive. In this case, the conjugacy class of an edge stabilizer of a one-edge $\mathbb{Z}$-splitting $T$ is uniquely determined, and we call it the core of $T$.
For a conjugacy class $\gamma$ of an infinite cyclic subgroup of $\mathbb{F}$, the intersection number of $\gamma$ and $T$, denoted by $i(\gamma,T)$, is defined as the minimal translation length of a nontrivial element of a subgroup representing $\gamma$ on $T$:
$i(\gamma,T)=\min\{\,d_{T}(x,wx)\mid 1\neq w\in c\in\gamma,\;x\in T\,\},$
where $d_{T}$ denotes the combinatorial metric on $T$. For example, if $\gamma$ is the core of $T$, then $i(\gamma,T)=0$.
The automorphism group $\operatorname{Aut}(\mathbb{F})$ contains the inner automorphism group $\operatorname{Inn}(\mathbb{F})$ as a normal subgroup, and the quotient
@@
is called the outer automorphism group of $\mathbb{F}$. Given a one-edge $\mathbb{Z}$-splitting $T$ of $\mathbb{F}$, one can associate a Dehn twist (or generalized transvection) $\delta_{\vec{T}}\in\operatorname{Out}(\mathbb{F})$, defined as an outer automorphism induced by twisting along $T$.
@@
The problem of constructing free subgroups of $\operatorname{Out}(\mathbb{F})$ has been extensively studied. Bestvina, Feighn, and Handel [1] proved that $\operatorname{Out}(\mathbb{F})$ satisfies the Tits alternative: every subgroup is either virtually solvable or contains a free subgroup of rank two. Kapovich and Lustig [13] showed that sufficiently large powers of two hyperbolic iwip elements generate either a virtually cyclic subgroup or a nonabelian free group. Uyanik [19] obtained a related result in the setting of hyperbolic extensions. Clay and Pettet [4] proved that if two one-edge $\mathbb{Z}$-splittings are filling, then sufficiently large powers of the associated Dehn twists generate a free group of rank two consisting of atoroidal fully irreducible elements except for the conjugacy classes of the twists themselves. Gültepe [9] provided another criterion under which powers of two Dehn twists generate a free subgroup of rank two.
A complementary development is due to Taylor [18], who proved that every right-angled Artin group embeds quasi-isometrically into some $\operatorname{Out}(\mathbb{F})$. He constructed a suitable collection of fully irreducible outer automorphisms supported on free factors, whose sufficiently large powers generate the desired RAAG. Since these automorphisms are fully irreducible, Dehn twists do not appear in his construction. Our work may therefore be viewed as a counterpart to Taylor’s theorem: we construct right-angled Artin subgroups of $\operatorname{Out}(\mathbb{F})$ generated instead by powers of Dehn twists.
Theorem 1.1 gives a broad and explicit method for constructing RAAG subgroups inside $\operatorname{Out}(\mathbb{F})$. However, the compatibility condition does not encompass all right-angled Artin phenomena arising from Dehn twists. For instance, $\operatorname{Out}(\mathbb{F}_{g})$ contains a direct product of 2g42g-4 free groups generated by Dehn twists (Horbez and Wade [11], and Bridson and Wade [3]) and a free-abelian subgroup of rank 2g32g-3 generated by Dehn twists (Feighn and Handel [7]), yet the associated one-edge $\mathbb{Z}$–splittings in these examples are not compatible. Thus, Theorem 1.1 captures a large and well-structured class of RAAGs generated by powers of Dehn twists, but it does not apply to every such instance found in $\operatorname{Out}(\mathbb{F}_{g})$.
We approach the proof of Theorem 1.1 by working in the mapping class group of the doubled handlebody
$\mathbb{M}=\mathbb{M}_{g}=\#_{g}(\mathbb{S}^{1}\times\mathbb{S}^{2}).$
The interaction between $\operatorname{Out}(\mathbb{F})$ and $\operatorname{MCG}(\mathbb{M})$ was first studied by Whitehead [20, 21] in his solution to the automorphic equivalence problem for $\mathbb{F}$. In particular, he proved that the natural map
$\operatorname{MCG}(\mathbb{M})\longrightarrow\operatorname{Out}(\mathbb{F})$
is surjective. Laudenbach [15] later established the short exact sequence
@@
Brendle, Broaddus, and Putman [2] subsequently proved that this sequence splits; in particular, there exists a section
$\operatorname{Out}(\mathbb{F})\longrightarrow\operatorname{MCG}(\mathbb{M}).$
In this setting, a Dehn twist may be interpreted as the isotopy class of a homeomorphism obtained by twisting $\mathbb{M}$ along an essential torus. Gültepe [9] likewise studied Dehn twists realized as twists along essential tori.
@@
One can formulate analogous criteria in the cases of two essential spheres or of an essential sphere together with an essential torus. In addition, one obtains a bigon criterion in this setting: an essential simple closed curve and an essential sphere (or an essential torus) intersect minimally if and only if they do not form a bigon. We note that Hatcher [10] introduced a normal form for sphere systems, and this normal form guarantees that any two sphere systems placed in it are in minimal position.
Meanwhile, Theorem 1.1 is proved by a ping–pong argument on the set of homotopy classes of essential simple closed curves. We employ a variation of the ping–pong method due to Koberda [14]. Careful control of intersection numbers within this framework yields the proof of Theorem 1.1.
In addition, we construct a compact topological space $\overline{(p\circ i_{*})(\mathcal{C})}$ on which $\operatorname{Out}(\mathbb{F})$ acts continuously. This space is precisely the closure of the set of points represented by homotopy classes of essential simple closed curves that represent $1$-free factors of $\mathbb{F}$. The definition of $\overline{(p\circ i_{*})(\mathcal{C})}$ is closely analogous to that of the projective space $\mathcal{PML}(S)$ of measured laminations on a surface $S$. In $\mathcal{PML}(S)$, the action of a Dehn twist is parabolic, and a Dehn twist in $\operatorname{Out}(\mathbb{F})$ exhibits the same dynamical behavior on $\overline{(p\circ i_{*})(\mathcal{C})}$, as described below.
@@
Let $T$ be a homotopy class of an essential torus with core $\gamma$. Let $\alpha$ be a homotopy class of an essential simple closed curve that represents a free factor. Then, we have $i(\alpha,T)>0$ if and only if $\delta_{\vec{T}}^{n}(\alpha)$ converges to $\gamma$ as $n\to\infty$.
1.1. Acknowledgments
The author thanks Sang-hyun Kim for many helpful and inspiring conversations during the preparation of this work. During the revision process, discussions with Michael Kopreski, Thomas Hill, George Shaji, and Brian Udall highlighted an error in a previous version of Lemma 3.6, leading to its improvement. The author is also grateful to Thomas Koberda for valuable comments that strengthened the manuscript. Thanks to Ignat Soroko, the author became aware of Whitehead’s work.
This paper is dedicated to my son, Jian, who arrived as this manuscript was taking shape.
@@ 2.1. Dehn twist automorphisms of free groups
An $\mathbb{F}$-tree $T$ is called a one-edge $\mathbb{Z}$-splitting if the action of $\mathbb{F}$ on the set of edges is transitive and a stabilizer of an edge is an infinite cyclic subgroup of $\mathbb{F}$. Let $\vec{E}T$ be the set of orientations of edges of $T$. Then there exists an orientation reversing $\vec{E}T\to\vec{E}T$, that is, $(\vec{e})$ is the reverse orientation of $\vec{e}$ for each $\vec{e}\in\vec{E}T$. Let $o:\vec{E}T\to\mathbb{F}$ be a map defined by
@@ 2.2. Dehn twists of doubled handlebodies
Let $\operatorname{MCG}(\mathbb{M})$ denote the mapping class group of $\mathbb{M}$, that is, the group of isotopy classes of orientation-preserving homeomorphisms. Laudenbach [15] shows the following short exact sequence holds:
@@
Brendle, Broaddus and Putman [2] discover that this sequence is splitting, that is, there exists a section $\operatorname{Out}(\mathbb{F})\hookrightarrow\operatorname{MCG}(\mathbb{M})$. It is therefore natural to refer to the image of a Dehn twist in $\operatorname{MCG}(\mathbb{M})$ simply as a Dehn twist. What kind of a homeomorphism is associated with a Dehn twist? To answer to this question, we need to dive into the geometry of $\mathbb{M}$.
@@
Note if two essential tori are homotopic, then they give the same Dehn twist as outer automorphisms. By Brendle, Broaddus and Putman [2], this outer automorphism can be considered as a mapping class. Note the mapping class group consists of isotopy classes of orientation-preserving homeomorphisms. So the following holds.
Proposition 2.7.
@@
We postpone the proof of Proposition 2.7 until after Lemma 3.6. Throughout this section, we will abuse notation by writing $T$ both for a homotopy class of an essential torus $t$ and for the associated dual tree of $t$. Similarly, we use $\vec{T}$ to denote either oriented version. With this convention, the Dehn twist along $\vec{T}$ will be denoted by $\delta_{\vec{T}}$.
@@ 3.1. Bigon, bihedron and minimal position
The following proof is adapted from Farb and Margalit [5].
@@
    Suppose there exist disks $D\subset t$ and $D^{\prime}\subset t^{\prime}$ such that $D\cup D^{\prime}$ forms a sphere bounding a $3$-ball. In this case, $t$ and $t^{\prime}$ are said to form a bihedron. By homotoping $t$ across this $3$-ball, one can eliminate the intersection circle $D\cap D^{\prime}$, thereby reducing the number of intersection circles between $t$ and $t^{\prime}$.
@@
    Suppose there are arcs $a\subset t$ and $a^{\prime}\subset t^{\prime}$ connecting two distinct intersection circles such that $a\cup a^{\prime}$ bounds a disk. Then $t$ and $t^{\prime}$ are said to form a bigon. By homotoping $t$ across this disk, one can reduce the number of intersection circles.
@@
Let $s$ and $t$ be essential spheres, essential tori, or one of each. Then, the number of intersection circles of $s$ and $t$ is exactly $i([s],[t])$ if and only if $s$ and $t$ form neither a bigon nor a bihedron.
@@
Choose two such circles, denoted by $C_{1}$ and $C_{2}$. Then there exist arcs $a\subset\hat{s}\times{0}$ and $b\subset S$, each connecting $C_{1}$ to $C_{2}$, such that the union $a\cup b$ bounds a disk in $\hat{s}\times[0,1]$. Since the union $H(a)\cup H(b)$ also bounds a disk, it follows that $s$ and $t$ form a bigon.
@@
By Laudenbach [15], two homotopic sphere systems are isotopic. Hatcher [10] show that two sphere systems can be in “normal form.” This is equivalent to minimal position. That is, two sphere systems are in normal form if and only if they are in minimal position. Refer to Horbez’s article [12].
The situation is not different if we consider an essential torus and a sphere system. Gültepe [8] defines the normal form of a torus with respect to a sphere system. This is equivalent to minimal position.
@@ 3.2. Intersection number and Dehn twist
Let $S$ be a homotopy class of either an essential sphere or an essential torus, and let $T$ be a homotopy class of an essential torus. If $i(S,\delta_{\vec{T}}^{k}(S))=0$ for all $k>0$, then we have $i(S,T)=0$.
@@
We investigate how essential spheres and essential tori evolve under iteration of a Dehn twist by examining the intersection between an essential surface and its image under a power of the twist. Let $T$ be a homotopy class of an essential torus with core $\gamma$. Let $k$ be a positive integer. Fix a homotopy class $S$ of an essential sphere with $i(S,T)>0$, and choose representatives $s\in S$ and $t\in T$ in minimal position.
For each component $c$ of $s\cap t$, exactly one of the following holds:
  1. (1)
    $c$ bounds a disk in $t$; or
  2. (2)
    $c$ represents a meridian of $t$.
The number of intersection circles of type (2) is equal to $i(\gamma,S)$. Consequently, the number of intersection circles of type (1) is $i(S,T)-i(\gamma,S)$. The local modification of a neighborhood of $c$ under the Dehn twist $\delta_{\vec{T}}^{k}$ depends essentially on whether $c$ is of type (1) or (2).
Figure 1. $c$ bounds a disk in $t$.
If (1) holds, that is, $c$ bounds a disk in $t$, then a neighborhood of $c$ in $s$ is stretched along the core $\gamma$ of $t$ under a positive power of the Dehn twist. As a result, this neighborhood intersects $s$ in
$k\,i(\gamma,S)+2$
circles. See Figure 1.
Figure 2. $c$ represents a meridian of $t$.
If (2) holds, that is, $c$ represents a meridian of $t$, then a neighborhood of $c$ in $s$ is stretched along the entire torus $t$ by a positive of the Dehn twist. In this case, the neighborhood intersects $s$ in
ki(S,T)1k\,i(S,T)-1
circles. See Figure 2.
Figure 3. The surfaces $s$ and $t$ intersect in three circles. In this configuration, there exists an essential torus homotopic to $\delta_{\vec{t}}(s)$ intersects $s$ in eight circles.
From the above local pictures, one obtains a corresponding global configuration. Figure 3 illustrates how these local modifications assemble into a global intersection pattern. Combining the contributions from each type of intersection circle, we obtain the following estimate: for every $k>0$,
$i(S,\delta_{\vec{T}}^{k}(S))$ $\geq(i(S,T)-i(\gamma,S))\,(k\,i(\gamma,S)+2)+i(\gamma,S)(k\,i(S,T)-1)$
$=k\,i(\gamma,S)\,(2\,i(S,T)-i(\gamma,S))+2\,i(S,T)-3\,i(\gamma,S).$
To justify this inequality, it suffices to verify that the resulting global configuration contains neither bigons nor bihedra. First, none of the local modifications described above produces a bihedron; consequently, no bihedron appears in the global picture. Next, suppose for contradiction that there exists a bigon between $s$ and $s_{k}$ where $s_{k}$ is a representative of $\delta_{\vec{T}}^{k}(S)$ locally described as the above. Then the arc of $s_{k}$ forming one side of the bigon can be decomposed into a concatenation of arcs, each of which lies entirely in either $s$ or $t$. It follows that there exists a bigon formed by $s$ and $t$. This contradicts the assumption that $s$ and $t$ are in minimal position. Therefore, no bigons occur between $s$ and $s_{k}$, and the inequality follows.
Thus $k\geq 2$ is a sufficient condition to guarantee that $i(S,\delta_{\vec{T}}^{k}(S))>0$. Equivalently, if $i(S,T)>0$ and $k\geq 2$, then
$i(S,\delta_{\vec{T}}^{k}(S))>0.$
This is the desired conclusion.
Suppose two homotopy classes $T$ and $T^{\prime}$ of essential tori have a positive intersection number, that is, $i(T,T^{\prime})>0$. Choose representatives $t\in T$ and $t^{\prime}\in T^{\prime}$ in minimal position. We say an essential simple closed curve of $t$ is a meridian if it bounds a disk in $\mathbb{M}$. Then each intersection circle $c$ of $t$ and $t^{\prime}$ satisfies one of the following:
  1. (1)
    $c$ bounds a disk in each of $t$ and $t^{\prime}$;
  2. (2)
    $c$ bounds a disk in $t$ and represents a meridian of $t^{\prime}$;
  3. (3)
    $c$ represents a meridian of $t$ and bounds a disk in $t^{\prime}$; or
  4. (4)
    $c$ represents a meridian of both $t$ and $t^{\prime}$.
The type of each intersection circle in the above list is invariant under homotopy. More precisely, for any other choice of representatives of $T$ and $T^{\prime}$ in minimal position, each of the above types of intersection circles occurring in $t\cap t^{\prime}$ also occurs for the new representatives.
Let us examine how the local picture near an intersection circle changes after applying a positive power of the Dehn twist $\delta_{\vec{T}}^{k}$ to $T^{\prime}$. Fix representatives $t\in T$ and $t^{\prime}\in T^{\prime}$ in minimal position, and let $c$ be a component of $t\cap t^{\prime}$.
Type (1). Suppose that $c$ bounds a disk in each of $t$ and $t^{\prime}$. Then there exist disks $D\subset t$ and $D^{\prime}\subset t^{\prime}$ with $\partial D=\partial D^{\prime}=c$ such that $D\cup D^{\prime}$ is an essential sphere; see the left-hand side of Figure 4. A representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ can be chosen so that, in a neighborhood of $c$, it has the local form shown on the right-hand side of Figure 4. In particular, this local modification creates at least two transverse intersection curves with $t^{\prime}$.
Figure 4. An intersection circle bounds a disk in each of $t$ and $t^{\prime}$.
Type (2). Suppose that $c$ bounds a disk in $t$ and represents a meridian of $t^{\prime}$. If $c$ appears as on the left-hand side of Figure 5, then a representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ has the local form shown on the right-hand side of Figure 5. In particular, in this neighborhood the surfaces $\delta_{\vec{T}}^{k}(t^{\prime})$ and $t^{\prime}$ intersect in at least two circles.
Figure 5. An intersection circle bounds a disk in $t$ and represents a meridian of $t^{\prime}$.
Type (3). Suppose that $c$ represents a meridian of $t$ and bounds a disk in $t^{\prime}$. Then $t$ and $t^{\prime}$ meet locally as on the left-hand side of Figure 6, and a representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ has the local form shown on the right-hand side. In this case the number of intersection circles of $t^{\prime}$ and $\delta_{\vec{T}}^{k}(t^{\prime})$ contained in a sufficiently small neighborhood of $c$ grows linearly with $k$. More precisely, this number is bounded below by $k\,i(T,T^{\prime})-1$.
Figure 6. An intersection circle represents a meridian of $t$ and bounds a disk in $t^{\prime}$.
Type (4). Suppose that $c$ represents a meridian of both $t$ and $t^{\prime}$. Then $t$ and $t^{\prime}$ meet locally as on the left-hand side of Figure 7, and a representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ has the local form shown on the right-hand side. In this case the number of intersection circles of $t^{\prime}$ and $\delta_{\vec{T}}^{k}(t^{\prime})$ contained in a sufficiently small neighborhood of $c$ equals $k\,i(T,T^{\prime})$.
Figure 7. An intersection circle represents a meridian of both $t$ and $t^{\prime}$.
We now estimate $i(\delta_{\vec{T}}^{k}(T^{\prime}),T^{\prime})$. For each $k>0$, construct a representative $t_{k}^{\prime\prime}$ of $\delta_{\vec{T}}^{k}(T^{\prime})$ by applying the local modifications described in (1)–(4) in a neighborhood of every intersection circle of $t\cap t^{\prime}$. In general, the pair $(t_{k}^{\prime\prime},t^{\prime})$ need not be in minimal position: the local construction may create bigons between $t_{k}^{\prime\prime}$ and $t^{\prime}$. However, the number of such bigons is uniformly bounded (independent of $k$), since bigons can only occur in finitely many prescribed neighborhoods determined by the original configuration of $t\cap t^{\prime}$. Therefore, after removing all bigons by homotopy (i.e. by repeatedly applying the bigon criterion), the total number of intersection circles decreases by at most a uniform constant.
On the other hand, the local models in Figures 6 and 7 produce at least $k\,i(T,T^{\prime})^{2}-i(T,T^{\prime})$ intersection circles between $t_{k}^{\prime\prime}$ and $t^{\prime}$ whenever there is at least one intersection circle of type (3) or (4). Consequently, for $k$ sufficiently large, the intersection number remains positive after bigon removal, that is,
$i(\delta_{\vec{T}}^{k}(T^{\prime}),T^{\prime})>0$
for all sufficiently large $k$. This is the desired statement. ∎
One of the implications from Lemma 3.6 is Proposition 2.7.
Proof of Proposition 2.7.
Lemma 2.6 implies that a homotopy class of an (oriented) essential torus determines an isotopy class of a Dehn twist. Thus, to prove Proposition 2.7, it suffices to show that if two Dehn twist homeomorphisms are isotopic, then their corresponding essential tori are homotopic.
Suppose that $t$ and $t^{\prime}$ are essential tori that are not homotopic, however, in minimal position. We claim that there exists an essential sphere $s$ such that either
$i([s],[t])=0<i([s],[t^{\prime}])\quad\text{or}\quad i([s],[t])>0=i([s],[t^{\prime}]).$
Let $c$ be a meridian of $t$. If $D$ is an embedded disk with $t\cap D=c=\partial D$, then one obtains an essential sphere $s_{D}$ as a boundary component of a regular neighborhood of $t\cup D$. If there exists such a disk $D$ with $i([s_{D}],[t^{\prime}])>0$, then $s_{D}$ satisfies the claim.
Otherwise, that is, if $i([s_{D}],[t^{\prime}])=0$ for all $D$, then there exists an essential sphere $s$ that intersects $t$ exactly at $c$. By the bigon–bihedron criterion, $s$ and $t$ are in minimal position. Since $t^{\prime}$ is disjoint from $s$, this sphere $s$ satisfies the claim. In conclusion, the claim holds.
By Lemma 3.6, there exists $k>0$ such that either
$i([s],[\delta_{\vec{t}}^{k}(s)])=0<i([s],[\delta_{\vec{t}^{\prime}}^{k}(s)])\quad\text{or}\quad i([s],[\delta_{\vec{t}}^{k}(s)])>0=i([s],[\delta_{\vec{t}^{\prime}}^{k}(s)]).$
Therefore, $\delta_{\vec{t}}$ is not isotopic to $\delta_{\vec{t}^{\prime}}$. ∎
If two homeomorphisms have disjoint supports, then they commute. Consequently, their mapping classes also commute. In other words, mapping classes that admit representatives with disjoint supports must commute. However, in general, commutativity does not imply the existence of representatives with disjoint supports. On surfaces, it is known that Dehn twists commute if and only if they admit representatives with disjoint supports (see [5, Fact 3.6]). The following theorem shows that the same phenomenon holds for doubled handlebodies.